OpenStudy (anonymous):

Use Long Division to simplify: 2 x squared minus 5x + 3 divided by 2x + 1

OpenStudy (anonymous):

\[\frac{2x^2-5x+3}{2x+1}=x-3+\frac{6}{2x+1}\]

OpenStudy (anonymous):

thank you